Task

How many grams of iron oxide can be produced with 56.3 grams of o2 and 57.8 grams of fe

Basic explanations needed
1
Expert's answer
2020-04-13T02:26:33-0400

Firstly, we need to check if there is one of the reagents (O2 or Fe) in excess (it means that some amount either of O2 or Fe might be not reacted). In order to do this we will calculate (according to chemical equation, see below) how many grams of O2 is needed to oxidize 57.8 grams of Fe:

4Fe + 3O2 = 2Fe2O3 (1)

m(O2) = (3*31.998 g/mol * 57.8 g)/4*55.847 g/mol = 24.84 g.

31.998 g/mol - Molar mass of Oxygen (calculated from Periodic table)

55.847 g/mol - Molar mass of Iron (Fe) (calculated from Periodic table)

Therefore to oxidize 57.8 g of Fe we need only 24.84 g of O2, what means that we have excess of O2 and 56.3 g - 24.84 g = 31.46 g of O2 will not react.

Secondly, we will calculate how many grams of Iron oxide (Iron (III) oxide) will form acording to reaction (1). So,

m(Fe2O3) = (2*159.69 g/mol*57.8 g)/4*55.847 g/mol = 82.64 g (of Fe2O3, iron oxide is formed)

159.69 g/mol - Molar mass of Fe2O3.

This amount also can be calculated using reacted mass of O2 (24.84 g):

m(Fe2O3) = (2*159.69 g/mol*24.84 g)/3*31.998 g/mol = 82.64 g (of Fe2O3, iron oxide is formed).

Answer: 82.64 g of iron oxide can be produced. (However, some amount of O2 (31.46 g) will not react)